### v3

# Rational Metareasoning for Large Language Models

Abstract
Recent approaches for leveraging large language models (LLMs) for reasoning often rely on additional inference-time computation to tackle complex tasks. However, as LLMs grow in size, the associated inference costs are becoming increasingly prohibitive. To address this, we introduce RaM (Rational Metareasoning): a cost-aware reasoning approach inspired by computational models of metareasoning in cognitive science. Our method trains LLMs to use intermediate reasoning steps selectively when they are likely to be beneficial. We first design a reward function that incorporates the Value of Computation by penalizing unnecessary reasoning, and then use it within an Expert Iteration framework to optimize the trade-off between performance and efficiency. Compared to few-shot chain-of-thought prompting and STaR, our method significantly reduces inference costs (23-45% fewer tokens generated across three models) while maintaining or increasing task performance across diverse datasets.
1 Introduction
Large language models (LLMs) rely on substantial computational power to handle complex problems (OpenAI et al., 2024; Chowdhery et al., 2022; de Vries, 2023). While initial studies mostly focused on the cost of training (Verdecchia et al., 2023), widespread use of LLMs has made inference-time costs an increasingly important factor. Moreover, there is a fundamental tension between inference cost and task performance: although chain-of-thought prompting (CoT; Wei et al., 2023; Kojima et al., 2023) and similar approaches improve task performance, they also substantially increase inference costs (Snell et al., 2024).
This trend has recently spiked with the development of Large Reasoning Models (OpenAI, 2024; DeepSeek-AI et al., 2025), instruction-tuned LLMs which are typically post-trained using reinforcement learning (RL) with outcome-based rewards to produce long chains of thought before each response. It is also worth noting that these approaches are not inherently adaptive: inference optimization (Wan et al., 2024) and existing CoT training methods often tend to raise or lower the inference cost on all queries, regardless of task complexity, or require the user to specify a budget up front (Anthropic et al., 2025).
In stark contrast to this static tradeoff, humans are able to adaptively allocate computational resources based on task difficulty (Kahneman, 2014; Russell, 1997; Lieder & Griffiths, 2017). In this work, we draw inspiration from rational metareasoning – literally, reasoning about reasoning – a concept originally from the artificial intelligence literature (Russell & Wefald, 1991) that has been used to explain how humans adaptively manage computational resources (Lieder & Griffiths, 2017; Lieder et al., 2018; Griffiths et al., 2019).
Building on this, we develop a novel reward function based on the Value of Computation (VOC; Russell & Wefald, 1991), which formalizes the trade-off between inference cost and task performance. We adopt an iterative reinforcement learning process inspired by the Expert Iteration algorithm (Anthony et al., 2017). In each iteration, we generate multiple reasoning chains for each question. These reasoning chains are ranked using the reward function, and the dataset is filtered to retain only the best reasoning chain for each question. The model is then fine-tuned using this filtered dataset. However, unlike previous applications of Expert Iteration to LLMs (Zelikman et al., 2022; Havrilla et al., 2024), which filter generated examples solely based on the correctness of the final answer, our method optimizes for both correctness and the cost of the reasoning process.
We evaluated the effectiveness of our solution across a diverse set of tasks, from scientific knowledge (ARC; Clark et al., 2018) to commonsense reasoning (CommonsenseQA; Talmor et al., 2019), mathematical problem solving (GSM8K; Cobbe et al., 2021), and logical deductive reasoning (ProofWriter; Tafjord et al., 2021). Additionally, we assess the out-of-domain generalization on MMLU (Hendrycks et al., 2021), a multitask benchmark. Our approach achieves a substantial reduction in generated tokens (35-42% compared to few-shot prompting, and 23-32% compared to STaR) while matching or increasing performance. Thus, we make the following contributions:
  1. 1.
    We employ rational metareasoning to optimize the tradeoff between inference cost and performance of LLMs on reasoning tasks.
  2. 2.
    We formalize a novel reward function inspired by the Value of Computation (VOC) and integrate it into LLM training.
  3. 3.
    We empirically demonstrate that rational metareasoning improves task performance at lower inference costs (23-42% fewer tokens on average) across various datasets and reasoning tasks.
2 Rational Metareasoning
Humans have limited time and cognitive resources (Griffiths et al., 2019; Griffiths, 2020). We face diverse challenges requiring different approaches: avoiding a sudden obstacle when driving needs quick, intuitive thinking, whereas selecting a retirement investment strategy requires slow, deliberate reasoning (Kahneman, 2014). Rational metareasoning (Russell & Wefald, 1991) captures this by suggesting agents should adapt their reasoning based on the problem at hand.
Intuitively, while reasoning solves a problem, metareasoning solves the problem of how to solve a problem: deciding which computations to perform while problem-solving. The essence of rational metareasoning is calculating the value of computation (VOC; Russell & Wefald, 1991) for each potential computation. The VOC balances the benefit of computation $c$ (the expected increase in the agent’s utility) against its cost (usually time or energy).
To formalize this, agents are assumed to have some internal belief state $b\in\mathcal{B}$, which determines their expectation about the value of each action $a\in\mathcal{A}$: $\mathbb{E}[U(a)|b]$. A rational agent would simply choose the highest-value action: $a^{*}=\text{argmax}_{a\in\mathcal{A}}{[U(a)|b]}$. However, this picture becomes more complex if an agent can perform computation to change their belief state before choosing an action. Each computation $c\in\mathcal{C}$ updates the agent’s belief to $b^{\prime}$ with probability $P(b^{\prime}|c)$, which in turn affects their beliefs about the value of actions, and incurs an associated cost ($\text{cost}(c)$). The VOC quantifies the value of performing computation $c$ given a starting belief state $b$,
$VOC(c,b)=\mathbb{E}_{P(b^{\prime}|c)}[\max_{a^{\prime}}\mathbb{E}[U(a^{\prime})|b^{\prime}]-\max_{a}\mathbb{E}[U(a)|b]]-\text{cost}(c).$ (1)
Thus, a meta-rational agent should pursue the computation $c^{*}$ with the highest VOC: $c^{*}=\text{argmax}_{c\in\mathcal{C}}VOC(c,b)$. If no computation has positive VOC, the agent should stop thinking and act in the world. Rational metareasoning can explain how humans allocate cognitive resources in various tasks (Lieder & Griffiths, 2017; Lieder et al., 2018; Callaway et al., 2018; 2021; 2022; Russek et al., 2022).
3 Rational Metareasoning with Large Language Models
To achieve an optimal balance between performance and efficiency, our approach introduces a new VOC-inspired reward function (Eq. 2) into Expert Iteration (Anthony et al., 2017; Zelikman et al., 2022), fine-tuning a LLM to produce reasoning chains adaptively.
Figure 1: RaM training. We iteratively generate reasoning chains using the current policy, score and filter them to approximate the optimal policy, and then finetune the base policy.
3.1 Reward modeling
Chain-of-thought prompting encourages LLMs to generate an intermediate output – a “chain of thought” – prior to producing the answer to a question (Wei et al., 2023; Kojima et al., 2023). We define the reward of a chain of thought as the difference between its utility and its cost,
$\mathcal{R}_{\pi}(x,y,z)=\mathcal{U}_{\pi}(z|x,y)-\mathcal{C}(z)$ (2)
where $x$ denotes the input for the task, $z$ represents the chain of thought, and $y$ is the target solution. The utility of the chain of thought is represented by $\mathcal{U}_{\pi}(z|x,y)$, and the cost of the intermediate computations is denoted by $\mathcal{C}(z)$. Equation 2 mirrors the VOC Equation 1: here, individual reasoning tokens correspond to intermediate computations $c$, making the reasoning chain $z$ a sequence of computations $c$, while the actions $a\in\mathcal{A}$ map to the potential outputs $y\in\mathcal{Y}$ of the language model. In the context of LLMs, utility quantifies the increase in the likelihood of generating the target sequence $y$ when the chain of thought $z$ is added to the input $x$, under the policy $\pi$:
$\mathcal{U}_{\pi}(z|x,y)=\log{\pi_{\theta}(y|z,x)}-\log{\pi_{\theta}(y|x)}.$ (3)
Specifically, $\pi_{\theta}(y|z,x)$ indicates the probability of generating the target sequence $y$ given both the chain of thought $z$ and the input $x$, while $\pi_{\theta}(y|x)$ denotes the probability of generating $y$ with only the input $x$. With respect to Equation 1, the language model’s initial belief about the value of actions or outputs is described by $\pi_{\theta}(y\mid x)$ (from Equation 3), whereas its final belief after generating a sequence of computations or tokens (a chain of thought $z$) is described by $\pi_{\theta}(y\mid z,x)$. The cost is directly proportional to the number of tokens in the chain of thought $l(z)$:
$\mathcal{C}(z)=\gamma\cdot l(z).$ (4)
The hyperparameter $\gamma$ scales the cost and utility to the same magnitude. A key benefit of this reward function is that it is parameterized by the same weights $\theta$ as the generative policy $\pi_{\theta}$, eliminating the need for an external reward model. This allows for direct estimation of the utility of a reasoning chain using the policy itself.
Finally, it’s worth noting that while standard reward models for RLHF bias the policy to produce longer reasoning chains (Singhal et al., 2024), thus necessitating appropriate corrections (Chen et al., 2024; Park et al., 2024) to correct for reward hacking, our reward model aims for and achieves the opposite effect.
3.2 Metareasoning Training
We demonstrate the effectiveness of our reward model using a variation of the Expert Iteration algorithm (EI, Anthony et al., 2017). EI is known for its sample efficiency and strong performance on reasoning tasks (Havrilla et al., 2024; Zelikman et al., 2022). As an example of an online reinforcement learning algorithm, EI involves both exploration and policy improvement phases, with the policy $\pi_{\theta}$ being updated using data from the exploration phase (as can be seen in Algorithm 1 and Figure 1).
Initially, in the exploration phase, we approximate the optimal policy $\hat{\pi}^{*}$ (whose computations maximize the VOC reward) by using rejection sampling on our student policy $\pi_{\theta}$. In particular, we begin with a pretrained language model $\pi_{\theta}$ and an initial dataset of problems $x$ along with their corresponding correct final answers $y$: $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{D}$. Following prior work in online RL (Tang et al., 2024), we utilize the model itself to generate the reasoning chains. Since we start with a pretrained model, we guide the model generation using few-shot prompting in the first iterations. Specifically, we concatenate a small set of examples, denoted as $\mathcal{P}$, each containing intermediate reasoning chains $z$, to each example in $\mathcal{D}$. During training and evaluation, we remove the few shot examples.
For each task $\tau_{i}=(x_{i},y_{i})$ in the original dataset $\mathcal{D}$, we generate $K$ reasoning chains: $\hat{\tau}_{i}=\{(x_{i},z_{k,i},y_{i})\}_{k=1}^{K}$. Then, we evaluate them using our reward function $\mathcal{R}_{\pi}$ (Section  3.1), and compute the advantage of each chain by subtracting the average reward for that question: $a_{i,k}=r_{i,k}-\frac{1}{K}\sum_{k^{\prime}=1}^{K}r_{i,k^{\prime}}$.
Using these advantage scores, we perform rejection sampling (by discarding reasoning chains with negative advantage) to construct the dataset of optimal trajectories $(\mathcal{D}_{1}^{*})$. Finally, we distill the selected rollouts into a policy $\pi_{1}$ via standard cross-entropy loss. This process can be iteratively repeated to refine the policy $\pi_{n}$ on the dataset $\mathcal{D}_{n}^{*}$. In closed-answer settings with verifiable answers, the reasoning chains that lead to incorrect answers can also be directly discarded, to produce a higher quality dataset.
Algorithm 1 Rational Metareasoning Training
Input $\pi$: a pretrained LLM; dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{D}$
1:$\pi_{0}\leftarrow\pi$triangleright Copy the original model
2:for $n$ in $1...N$ do
3:     for $k$ in $1...K$ do
4:         $({z}_{i,k},{y}_{i,k})\leftarrow\pi_{n-1}(x_{i})\quad\forall i\in[1,D]$triangleright Sample reasoning chains
5:     end for
6:     $r_{i,k}\leftarrow\mathcal{R}_{\pi_{n-1}}(x_{i},y_{i},z_{i,k})\quad\forall i,k(i\in[1,D],k\in[1,K])$triangleright Compute rewards
7:     $a_{i,k}\leftarrow r_{i,k}-\frac{1}{K}\sum_{k^{\prime}=1}^{K}r_{i,k^{\prime}}\quad\forall i,k(i\in[1,D],k\in[1,K])$triangleright Compute advantages
8:     $\hat{Z}_{i}\leftarrow\{\,z_{i,k}\mid a_{i,k}>0\,\}$triangleright Select best reasoning chains
9:     $\mathcal{D}_{n}^{*}\leftarrow\{\,(x_{i},\hat{z}_{i,k},y_{i})\mid\hat{z}_{i,k}\in\hat{Z}_{i},\,i\in[1,D]\,\}$triangleright Create the optimal dataset
10:     $\pi_{n}\leftarrow\text{train}(\pi,\mathcal{D}_{n}^{*})$triangleright Finetune the original model on the optimal solutions
11:end for
4 Experiments
4.1 Datasets
To characterize the generality of our method, we applied it to a diverse range of datasets and reasoning tasks. We constructed our training set by combining the training sets from the following datasets into one dataset $\mathcal{D}$ and then evaluated the model on all corresponding public test sets $\mathcal{T}$. To facilitate comparison with baselines that use the correctness of the answer as a reward, we limited the training to datasets with verifiable answers, although this is not strictly required by our method. We used four datasets:
ARC (Clark et al., 2018). The AI2 Reasoning Challenge (ARC) dataset comprises grade-school science questions, designed to evaluate the capability to apply scientific knowledge.
CommonsenseQA (Talmor et al., 2019). This dataset is centered on commonsense question answering. It leverages implicit human knowledge and testing everyday reasoning.
GSM8K (Cobbe et al., 2021). This dataset includes a variety of linguistically diverse grade-school math word problems. It assesses proficiency in solving mathematical problems that require comprehension and application of arithmetic reasoning.
ProofWriter (Tafjord et al., 2021). This dataset assesses logical deductive reasoning by asking the model to determine if a conclusion follows from premises presented in natural language.
These datasets have very different train split sizes. To ensure fairness and balance between the datasets, and to manage computational costs, we composed our training mixture by sampling 1024 random samples from each of the training sets (for a total of 4096 samples). We then evaluated the model on the public test set of each dataset. To further assess the generalization of our approach, we conducted out-of-distribution testing on MMLU-CF (Zhao et al., 2024), a contamination free and more challenging variant of the original MMLU dataset (Hendrycks et al., 2021). This dataset consists of 10,000 multiple-choice questions from various branches of knowledge.
4.2 Baselines
We illustrate the advantages of our model by comparing its performance to two types of prompting strategies: Direct prompting, where the model is required to provide an immediate answer, and Chain of Thought prompting (CoT), where the model is encouraged to reason through the problem step-by-step before arriving at a solution. Since we are using pretrained models which are not specifically trained for instruction following, we provide five few-shot examples for each task from the unused portion of the training dataset. These are the same examples that are used to guide the model training during the first iteration. In addition to these prompting methods, we adopt as a finetuning baseline the most common reasoning bootstrapping method, STaR (Self-Taught Reasoner; Zelikman et al. 2022), which also uses a variation of Expert Iteration. In particular, this method was the first to adopt outcome-based rewards to enable models to self-improve: given a set of problems and solutions, it trains the model iteratively on reasoning chains that lead to the correct solution.
As an upper bound, we also tested the instruction-tuned version of the models, which have been fine-tuned on a massive amount of annotated chain-of-thought data, followed by RLHF (Dubey et al., 2024). Finally, while CoT prompting may not yield optimal trajectories, more advanced methods (Yao et al., 2023a; Zheng et al., 2024; Madaan et al., 2023) often increase sequence lengths substantially, and are therefore less efficient. We focus on CoT for its simplicity – reducing reasoning tokens while gaining performance for CoT suggests similar gains for more complex approaches.
4.3 Training Details
For our experiments, we use Meta Llama-3.2-3B and Llama-3.1-8B (Dubey et al., 2024) as the pretrained base models. We have chosen $\gamma=0.1$ to align the distributions of costs and rewards more closely, although we have found our method to be robust to small variations in the choice of this hyperparameter. We sample $K=4$ reasoning chains for each question, using a temperature $t$ of 0.5 and a $top_{p}$ value of 0.9. These parameters are chosen to balance exploration and exploitation, allowing us to generate diverse yet relevant reasoning chains.
For greater efficiency during training, instead of using the entire dataset from the start, we begin by sampling a dataset $\mathcal{D}_{n}$ of size $T=1024$ from the union of the four training datasets $\mathcal{D}$ described in 4.1. We then progressively increase the dataset size by $T$ with each subsequent iteration (until it reaches the size of $\mathcal{D}$) allowing the model to encounter new examples gradually. In the self-supervised fine-tuning step, we use a batch size of 16 and a learning rate of 1e-5. We execute five iterations of the training algorithm. We believe that further improvements are possible through a more comprehensive hyperparameter search; however, due to computational constraints, we leave this for future work. Finally, we evaluate all models using greedy decoding to ensure consistent and deterministic output generation. We use pattern matching techniques to extract the answers; an exact match between the generated answer and the ground truth is considered correct.
5 Results
5.1 Performance vs Cost
We first evaluate our approach against baselines (Sec. 4.2) across several datasets (Sec. 4.1). Our key criteria are performance (measured by rescaled accuracy) and cost (measured by the number of input and output tokens). Our experiments confirm that across all models and datasets, our training approach reduces cost while matching or improving performance (see Table 1 for results averaged across datasets; Tables 4 and 5 for per-dataset results; and Appendix C for example reasoning chains). Fig. 2 shows these results for all models.
Figure 2: Cost and performance. Accuracy is plotted against the output tokens. Our method, RaM, eliminates the need for few-shot prompting (reducing input tokens) and trains the model to use fewer reasoning tokens than STaR.
Model Method Accuracy (%) $\uparrow$ # Input Tokens $\downarrow$ # Output Tokens $\downarrow$
Llama-3.2-3B Direct Few-Shot 33.5 ($\pm$ 1.4) 291.7 ($\pm$ 0.9) 0.0 ($\pm$ 0.0)
CoT Few-Shot 34.8 ($\pm$ 1.8) 1080.5 ($\pm$ 0.9) 261.9 ($\pm$ 4.8)
STaR 47.4 ($\pm$ 2.0) 73.2 ($\pm$ 0.9) 222.7 ($\pm$ 3.5)
RaM 50.1 ($\pm$ 1.8) 73.2 ($\pm$ 0.8) 170.8 ($\pm$ 2.9)
Llama-3.2-3B-Instruct CoT Prompting 53.2 ($\pm$ 1.8) 157.2 ($\pm$ 0.9) 259.1 ($\pm$ 3.4)
Llama-3.1-8B Direct Few-Shot 45.8 ($\pm$ 1.5) 291.7 ($\pm$ 0.9) 0.0 ($\pm$ 0.0)
CoT Few-Shot 56.4 ($\pm$ 1.9) 1080.5 ($\pm$ 0.9) 253.6 ($\pm$ 4.2)
STaR 61.9 ($\pm$ 1.6) 73.2 ($\pm$ 0.9) 215.1 ($\pm$ 3.3)
RaM 63.5 ($\pm$ 1.7) 73.2 ($\pm$ 0.9) 146.2 ($\pm$ 3.3)
Llama-3.1-8B-Instruct CoT Prompting 60.9 ($\pm$ 1.7) 157.2 ($\pm$ 0.9) 269.9 ($\pm$ 3.8)
Table 1: Comparison of different methods based on rescaled accuracy and length metrics, averaged across datasets (means with 95% confidence intervals; bold indicates best performing approaches with overlapping 95% intervals). RaM achieves better performance while using significantly fewer input and output tokens compared to STaR or CoT Few-Shot.
We first consider the performance of our baselines: CoT Few-Shot prompting uses a large number of input and output tokens, but yields reasonable performance. Direct Few-Shot prompting uses fewer input (and far fewer output) tokens, but yields poor performance on reasoning-intensive datasets (GSM8K and Proofwriter, see Table 4). STaR improves on these approaches, using significantly fewer tokens while achieving comparable or superior performance.
RaM further improves the cost-performance tradeoff by generating 23-32% fewer tokens on average compared to STaR, generally with higher accuracy. Interestingly, our findings align with the conclusions of a meta-analysis of over 100 papers (Sprague et al., 2024), in finding that explicit chains help mainly on maths or logic tasks and bring marginal gains elsewhere.
Finally, it is worth noting that while the performance of the instruction-tuned model (CoT Instruct) is generally higher than that of the pretrained model, the length of the reasoning chains was generally much higher and not meaningfully adaptive to task complexity.
5.2 Adaptive computation
Section 5.1 demonstrates that our method reduces computational costs on average. But does it actually teach models to reason adaptively (by adjusting reasoning to match task complexity), or just to reason less? To address this question, we first divided our test set $\mathcal{T}$ based on whether or not reasoning was needed to solve the task. In particular, we split the data based on whether Direct Few-Shot obtained the correct answer (“easy split”) or not (“hard split”). Adaptive methods should use less computation to solve the easy problems. We can empirically compare the results across methods for these two data splits.
As shown in Table 2, all models and methods are able to differentiate between hard and easy problems, generating fewer tokens on easier problems. The instruction-tuned model (CoT Instruct) did not effectively adapt response lengths to align with task complexity in our experiments. While STaR seems to improve this ratio, RaM increases the difference in reasoning between hard and easy problems, achieving a length reduction of up to 50.3% on Llama-3.2-3B. This indicates that RaM trains models to reason adaptively, helping them recognize when detailed reasoning is necessary and when a shorter response is sufficient.
5.3 Generalization
We assess the out-of-distribution generalization of the length reduction of RaM with respect to STaR using the public test set of the MMLU-CF benchmark (Hendrycks et al., 2021, see Section 4.1). As shown in Table 3, RaM achieves similar performance while generating 28% to 36% fewer tokens than STaR.
Looking at subdomains, Fig. 3 shows the ratio between the output length of RaM with respect to STaR. We can see that subjects that require multi-step reasoning (for example, physics or engineering) exhibit smaller reductions because they need detailed intermediate steps. In contrast, tasks focused primarily on recalling factual information tend to show a more pronounced length reduction.
Model Method Hard Split Easy Split Length Reduction (%) $\uparrow$
Llama-3.2-3B CoT Few-Shot 210.0 ($\pm$ 3.0) 170.0 ($\pm$ 3.0) 19.0
STaR 195.0 ($\pm$ 5.0) 128.0 ($\pm$ 1.0) 34.4
RaM 145.0 ($\pm$ 3.0) 72.0 ($\pm$ 2.0) 50.3
Llama-3.2-3B-Instruct CoT Prompting 222.0 ($\pm$ 4.0) 221.0 ($\pm$ 3.0) 0.5
Llama-3.1-8B CoT Few-Shot 197.5 ($\pm$ 3.5) 151.0 ($\pm$ 1.0) 23.5
STaR 202.0 ($\pm$ 5.0) 129.5 ($\pm$ 2.5) 35.9
RaM 139.0 ($\pm$ 3.0) 89.0 ($\pm$ 1.0) 36.0
Llama-3.1-8B-Instruct CoT Prompting 228.0 ($\pm$ 6.0) 226.0 ($\pm$ 2.0) 0.9
Table 2: Length of generated reasoning chains across models and methods (mean number of tokens with 95% confidence intervals). Adaptive methods should maximize the difference in length between the two distributions. RaM reduces the overall length and increases the difference in the length distribution between the Hard and Easy splits (Length Reduction), demonstrating an improvement in the ability to adapt reasoning length to task complexity.
Model Method Accuracy (%) $\uparrow$ # Input Tokens $\downarrow$ # Output Tokens $\downarrow$
Llama-3.2-3B STaR 27.0 ($\pm$ 2.4) 75.1 ($\pm$ 0.9) 150.6 ($\pm$ 4.1)
RaM 29.5 ($\pm$ 2.5) 75.1 ($\pm$ 0.9) 96.3 ($\pm$ 3.7)
Llama-3.1-8B STaR 40.2 ($\pm$ 2.4) 75.1 ($\pm$ 0.8) 153.6 ($\pm$ 4.0)
RaM 36.4 ($\pm$ 2.4) 75.1 ($\pm$ 0.9) 110.6 ($\pm$ 3.8)
Table 3: Comparison of STaR and RaM based on rescaled accuracy and length metrics in an out-of-distribution setting on the MMLU-CF benchmark. We report the mean with 95% confidence intervals. RaM has similar performance but generates fewer output tokens.
Figure 3: Ratio of output length reduction. RaM ’s output length relative to STaR’s, for Llama-3.2-3B. Tasks requiring heavier reasoning tend to exhibit a lower length reduction ratio, whereas those relying more on knowledge display a more pronounced length reduction.
6 Related Work
6.1 Reducing inference costs
The rising cost of deploying large language models (LLMs) has driven efforts to reduce inference costs. Techniques such as Speculative Decoding (Leviathan et al., 2023) and Medusa (Cai et al., 2024) improve efficiency through parallelization, while Mixture of Experts (Jacobs et al., 1991; Zhou et al., 2022) activates only a subset of LLM parameters during decoding. Though effective, these methods require significant architectural changes and don’t adapt computation based on task difficulty. Other approaches have developed neural architectures that enable adaptive computation (Graves, 2017; Banino et al., 2021; Dehghani et al., 2019; Mohtashami et al., 2023; Schuster et al., 2022) but involve new architectures or training methods. In contrast, our approach uses existing architectures and pretrained models, modifying only the fine-tuning process. More similar to our approach, model routing (Ong et al., 2024; Jiang et al., 2024) optimizes resource utilization based on query complexity by routing easier queries to smaller models and harder queries to larger ones. However, this necessitates multiple models and a router, while our approach trains a single model to adaptively adjust its own outputs to match task complexity.
Finally, the concurrent “long-to-short” literature deliberately compresses reasoning paths while preserving accuracy. (Arora & Zanette, 2025) shows that adding a length penalty to verifiable rewards can shorten the reasoning length of Large Reasoning Models in mathematical settings. Here, we focus on bootstrapping the reasoning process, teaching pretrained language models to reason efficiently from scratch by developing a reward that is agnostic to the answer format. CoT-Valve (Ma et al., 2025) and Claude 3.7 (Anthropic, 2025) enable external control over the length of the reasoning chains (and therefore the computation budget), allowing a model to produce both long and short reasoning chains. Unlike these methods, our method enables the model to independently adapt the reasoning length according to the difficulty of the task. Concurrent work (Wu et al., 2025; Luo et al., 2025) also shows that reasoning length compression can also be achieved by merging LLMs and Large Reasoning Models. Similarly, Synergy-of-Thoughts (Shang et al., 2024) employed models of different sizes to adapt to the complexity of the task at hand. Our approach differs from all of the above on two fronts. First, it keeps a single set of parameters: there is no external controller, merged checkpoint, or auxiliary model. Second, adaptivity is learned through a principled Value-of-Computation reward, not by hard length constraints or heuristic escalation.
6.2 Reasoning in LLMs
Techniques such as Chain of Thought (CoT) and related methodologies (Wei et al., 2023; Yao et al., 2023a; Madaan et al., 2023; Zheng et al., 2024) have proven effective at enhancing LLM performance across a wide range of tasks. CoT boosts LLMs’ performance on complex reasoning by guiding them through a series of intermediate reasoning steps, increasing inference costs to improve task performance. This method can be implemented through in-context learning (Wei et al., 2023), prompting (Kojima et al., 2023), or training (Li et al., 2023). The benefits of CoT can be attributed to both a greater computation depth (Goyal et al., 2024; Pfau et al., 2024) and the semantic values of the thought tokens, which function as intermediate variables in the computation of the answer (Prystawski et al., 2023). However, recent studies have raised concerns regarding the meaningfulness of such reasoning chains in reaching the target solution, and whether models effectively utilize them to solve tasks (Turpin et al., 2023; Paul et al., 2024; Sprague et al., 2024). We further demonstrate that standard prompting and training methods fail to teach the model to use CoT purposefully, resulting in inefficient inference. Moreover, reasoning in situations where it is unnecessary may also harm model performance Liu et al. (2024).
Reasoning can also be used to bootstrap language models. Self-improving techniques (Huang et al., 2022; Zelikman et al., 2022) consist of generating reasoning chain-augmented answers for unlabeled questions and fine-tuning the LLM using those self-generated solutions as target outputs. Similar techniques, such as ReST (Gulcehre et al., 2023), can also be used to better align LLMs with human preferences and needs. Our approach builds on these techniques to optimize the inference cost of reasoning in addition to task performance.
Another related work, Quiet-STaR (Zelikman et al., 2024), has improved LLM reasoning performance by modifying pretraining to generate intermediate reasoning sequences between tokens. While effective for downstream tasks, it increases computational costs by generating reasoning chains at every step, even when unnecessary. More recently, chat models that “think before answering” have been developed (OpenAI, 2024; Anthropic et al., 2025; DeepSeek-AI et al., 2025), using inference-time computation to enhance their outputs. Although these models outperform others, they expend more computational resources even when it may not be necessary. Our method could be incorporated into their training process to help the model determine when this additional computation is genuinely beneficial.
7 Limitations and Future Work
While our experiments show that rational metareasoning can successfully reduce inference costs in large language models, some limitations remain. First, our approach has only been tested on well-established datasets in science, commonsense reasoning, logic, and mathematics. Its effectiveness in other contexts has yet to be demonstrated. One particularly relevant example is the agentic setting, where LLMs act autonomously in complex digital environments (Yao et al., 2023b; Schick et al., 2023). Adapting our method to this context would require incorporating the cost of tool use (e.g., API calls) into the reward function to encourage the model to minimize unnecessary resource consumption. Second, while the VoC reward function is highly flexible and can be readily combined with various learning algorithms, our work focused on Expert Iteration (Havrilla et al., 2024) due to its sample efficiency for reasoning tasks. Extending the VoC reward framework to other training approaches like PPO or DPO remains an open research area. Evaluating whether VoC-based rewards maintain or even improve performance in these settings would further clarify the robustness and generalizability of our approach.
8 Conclusion
We introduced a cognitively inspired reward function grounded in rational metareasoning, aimed at optimizing LLM inference by balancing performance and computational cost. Empirically, we showed that this approach substantially reduces the number of generated tokens and input context length, while consistently improving task accuracy across a diverse range of benchmarks. Most excitingly, our work demonstrates how cognitively-inspired reward functions can endow LLMs with desirable inference-time properties, opening a broad avenue of future work. Given its flexibility, this method could be integrated into instruction tuning to potentially enhance performance, even in scenarios where verifying the correctness of answers is challenging. Since the utility measure within the reward function can be tailored to prioritize any desired, measurable property, this approach offers the potential to guide models toward achieving these enhanced qualities while still benefiting from the reduced computational costs.
Appendix A Complete Results Tables
In this section, we present the comprehensive results tables.
Model Method ARC Commonsenseqa GSM8K Proofwriter
Llama-3.2-3B Direct Few Shot 70.1 ($\pm$ 1.8) 56.2 ($\pm$ 3.2) 7.6 ($\pm$ 1.6) 0.0 ($\pm$ 1.3)
Cot Few Shot 63.8 ($\pm$ 1.9) 43.3 ($\pm$ 3.3) 27.7 ($\pm$ 2.4) 4.2 ($\pm$ 4.5)
Direct SFT 82.8 ($\pm$ 1.2) 66.0 ($\pm$ 2.6) 41.2 ($\pm$ 2.8) 36.3 ($\pm$ 3.1)
STaR 66.6 ($\pm$ 1.9) 52.5 ($\pm$ 3.5) 26.8 ($\pm$ 2.5) 43.6 ($\pm$ 4.5)
RaM 71.2 ($\pm$ 1.7) 55.8 ($\pm$ 3.3) 32.3 ($\pm$ 2.5) 41.3 ($\pm$ 4.4)
Llama-3.2-3B-Instruct CoT Prompting 75.9 ($\pm$ 1.7) 58.1 ($\pm$ 3.2) 71.6 ($\pm$ 2.2) 7.3 ($\pm$ 4.4)
Llama-3.1-8B Direct Few Shot 82.6 ($\pm$ 1.5) 66.4 ($\pm$ 3.0) 14.9 ($\pm$ 2.0) 19.4 ($\pm$ 4.4)
Cot Few Shot 80.9 ($\pm$ 1.5) 60.4 ($\pm$ 3.2) 52.8 ($\pm$ 2.6) 32.9 ($\pm$ 4.2)
STaR 82.0 ($\pm$ 1.6) 64.0 ($\pm$ 3.0) 45.9 ($\pm$ 2.8) 55.8 ($\pm$ 4.1)
Metareasoning 80.3 ($\pm$ 1.5) 67.2 ($\pm$ 3.1) 53.7 ($\pm$ 2.5) 52.7 ($\pm$ 4.0)
Llama-3.1-8B-Instruct CoT Prompting 85.4 ($\pm$ 1.4) 66.6 ($\pm$ 3.0) 80.5 ($\pm$ 2.0) 11.2 ($\pm$ 4.3)
Table 4: Comparison of accuracy of different methods across datasets and models. We report the mean with 95% confidence scores.
Model Method ARC Commonsenseqa GSM8K Proofwriter
Llama-3.2-3B Direct Few Shot 0.0 ($\pm$ 0.0) 0.0 ($\pm$ 0.0) 0.0 ($\pm$ 0.0) 0.0 ($\pm$ 0.0)
Cot Few Shot 169.6 ($\pm$ 2.3) 182.0 ($\pm$ 9.7) 236.9 ($\pm$ 8.0) 459.2 ($\pm$ 14.1)
STaR 121.0 ($\pm$ 2.2) 130.0 ($\pm$ 1.4) 259.7 ($\pm$ 9.0) 380.0 ($\pm$ 7.2)
RaM 77.4 ($\pm$ 1.5) 91.3 ($\pm$ 2.6) 184.1 ($\pm$ 7.8) 330.4 ($\pm$ 7.4)
Llama-3.2-3B-Instruct CoT Prompting 230.8 ($\pm$ 2.2) 189.1 ($\pm$ 3.7) 192.1 ($\pm$ 4.6) 424.6 ($\pm$ 11.0)
Meta-Llama-3-8B Direct Few Shot 0.0 ($\pm$ 0.0) 0.0 ($\pm$ 0.0) 0.0 ($\pm$ 0.0) 0.0 ($\pm$ 0.0)
Cot Few Shot 147.9 ($\pm$ 1.5) 153.6 ($\pm$ 6.8) 232.7 ($\pm$ 8.6) 472.1 ($\pm$ 16.8)
STaR 119.6 ($\pm$ 2.0) 143.6 ($\pm$ 2.3) 256.9 ($\pm$ 7.5) 340.4 ($\pm$ 9.6)
Metareasoning 91.8 ($\pm$ 1.4) 93.1 ($\pm$ 3.2) 171.4 ($\pm$ 5.4) 228.7 ($\pm$ 10.1)
Llama-3.1-8B-Instruct CoT Prompting 233.0 ($\pm$ 2.2) 200.7 ($\pm$ 4.3) 210.8 ($\pm$ 5.0) 435.1 ($\pm$ 11.7)
Table 5: Comparison of mean output length of different methods across datasets and models. We report the mean with 95% confidence scores.
Appendix B CoT Prompting
For the “Instruct” versions of our models we use the following prompt:
Answer the following question, thinking step by step to get to the answer. You can think however long you need, but answer as soon as you’re ready. Use the minimum number of steps to get to the answer. Once you’re finished thinking, you must end your response with ’The answer is [X]’, where [X] is the final answer to the question.
Appendix C Qualitative examples
Below we show one example of question, reasoning chain and answer per dataset, generated with each method: CoT Few Shots (blue), STaR (yellow) and RaM (green). The following examples were generated by Llama-3.2-3B (Dubey et al., 2024). The reasoning chains shown here serve as intermediate computations that guide the model to its final answers. While not intended as direct explanations, the examples remain interpretable and demonstrate the effectiveness of the respective methods.
Figure 4: Qualitative examples of reasoning processes for ARC dataset
Figure 5: Qualitative examples of reasoning processes for CommonsenseQA dataset
Figure 6: Qualitative examples of reasoning processes for GSM8k dataset
Figure 7: Qualitative examples of reasoning processes for Proofwriter dataset